OpenStudy (anonymous):

Check my answer? Will give medal! On April 11, 2012, two earthquakes were measured off the northwest coast of Sumatra. The first had a magnitude of 8.6. The second had a magnitude of 8.2. By what approximate factor was the intensity of the first earthquake greater than the intensity of the second earthquake?

OpenStudy (anonymous):

\[M=\log(\frac{ l }{ l _{0} })\] M = the magnitude of an earthquake I = the intensity of an earthquake I0 = the smallest seismic activity that can be measured

OpenStudy (anonymous):

0.40 1.02 1.05 2.51

OpenStudy (anonymous):

I believe it's 2.51, but I can't really word how exactly I got the answer so that my process makes sense. If anyone can help me out with an actual explanation and check the answer I've given, I would greatly appreciate it.

OpenStudy (perl):

hello, would you like help?

OpenStudy (anonymous):

Sure. Why not. :)

OpenStudy (perl):

On April 11, 2012, two earthquakes were measured off the northwest coast of Sumatra. The first had a magnitude of 8.6. The second had a magnitude of 8.2. By what approximate factor was the intensity of the first earthquake greater than the intensity of the second earthquake? Let M1 = magnitude of first earthquake M2 = magnitude of second earthquake I1 = intensity of first earthquake I2 = intensity of second earthquake Io = intensity of the smallest seismic activity that can be measured which is unknown to us, but which is a constant. M1 = log ( I1 / Io) M2 = log ( I2 / Io)

OpenStudy (anonymous):

I've kinda forgotten how to do this. How would you calculate the intensity to compare the two?

OpenStudy (perl):

you can substitute and use properties of logs

OpenStudy (perl):

We are given that M1 = 8.6 M2 = 8.2 substitute 8.6 = log( I1 / Io) 8.2 = log (I2 / Io) raise both sides to the power of 10 10^8.6 = 10^log(I1 / Io) 10^8.2 = 10^log(I2 / I0 )

OpenStudy (perl):

now 10^log(anything) = anything they cancel

OpenStudy (anonymous):

10^log(I1 / Io) would now just be I1 / Io and 10^log(I2 / Io) would be I2 / Io

OpenStudy (perl):

8.6 = log( I1 / Io) 8.2 = log (I2 / Io) raise both sides to the power of 10 10^8.6 = 10^log(I1 / Io) 10^8.2 = 10^log(I2 / Io) 10^8.6 = I1 /Io 10^8.2 = I2 / Io multiply both sides by Io 10^8.6 * Io = I1 10^8.2 * Io = I2 I2 / I1 = (10^8.6 * Io) / (10^8.2 * Io)

OpenStudy (perl):

now on the right side, the Io's cancel . remember its a constant

OpenStudy (perl):

I2 / I1 = (10^8.6 * Io) / (10^8.2 * Io) = 10^8.6 / 10^8.2 = 10 ^ (8.6 - 8.2) = 10^.4 = 2.51

OpenStudy (perl):

the reason why i divided I2 / I1, was because the question asked by what approximate factor is the intensity of earthquake 1 greater than earthquake 2

OpenStudy (anonymous):

Wow! So I was right? I don't even remember how I got it.

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (perl):

:)

OpenStudy (perl):

now you can close this month long question :)

OpenStudy (perl):

oh you did